OpenStudy (anonymous):

What is the fourth term of the expansion of the binomial (2x + 5)5

OpenStudy (anonymous):

In the expansion of \[\left( x+a \right)^n\] (r+1)th term is \[T _{r+1}=C _{r}^{n}x ^{n-r}a ^{r}\] put x=2x,a=5 n=5 and calculate.

OpenStudy (anonymous):

r+1=4 r=4-1=3